OpenStudy (anonymous):

solve the ODE: x''+5x'+4x=8

myininaya (myininaya):

different process for this one

myininaya (myininaya):

first lets solve x''+5x'+4x=0 r^2+5r+4=0 (r+1)(r+4)=0 r=-1, r=-4 so we have the homogeneous solution is \[y=c_1e^{-1t}+c_2e^{-4t}\]

myininaya (myininaya):

oops that y is suppose to be x

OpenStudy (anonymous):

and then if I have IB conditions I can just plug in and solve for C_1 and C_2?

OpenStudy (anonymous):

and if one IB is x'(0)=0, I will also need to take the derivative of homogeneous solution?

myininaya (myininaya):

we can check this \[x'=-c_1e^{-1t}-4c_2e^{-4t}=> x''=c_1e^{-1t}+16c_2e^{-4t}\] \[(c_1e^{-1t}+16c_2e^{-4t})+5(-c_1e^{-1t}-4c_2e^{-4t})+4(c_1e^{-1t}+c_2e^{-4t})\] \[=(1-5+4)e^{-1t}+(16-20+4)e^{-4t}=0e^{-1t}+0e^{-4t}=0\] so anyways this is only part of our solution \[x=c_1e^{-1t}+c_2e^{-4t}+t_p\] so we need to find the non-homogeneous solution we need to look back as x''+5x'+4x=8

myininaya (myininaya):

joe you wanna finish it? lol

OpenStudy (anonymous):

i cant remember the method to finding a particular solution lol >.< other than eye-balling it and saying x = 2 is a particular solution.

myininaya (myininaya):

we need to find that t_p part i think it is called the particular solution

myininaya (myininaya):

lol

myininaya (myininaya):

yes x=2 works

OpenStudy (anonymous):

lolol, theres a set way of finding it, i just cant remember, ima look it up.

OpenStudy (anonymous):

...unless phi knows it and is going to post it :)

myininaya (myininaya):

x=A+Bt+Ct^2 or something like that

myininaya (myininaya):

and then we find x' and x'' and find what A B and C are

myininaya (myininaya):

to make the equation true

myininaya (myininaya):

lets see what phi says :)

OpenStudy (phi):

You can find the particular solution by assuming a linear combination of the RHS and all its derivatives. This only works for things with a finite number of derivatives. But a constant works. Assume x= A. x' =0 and x''=0. Plug in and solve for A the answer is 2

myininaya (myininaya):

\[x=c_1e^{-1t}+c_2e^{-4t}+2\]

myininaya (myininaya):

so you said you had boundary conditions?

myininaya (myininaya):

no! stupid thing was going so slow so i clicked too many times and it deleted what i wrote :(

OpenStudy (anonymous):

>.< lol

OpenStudy (anonymous):

x'(0)=0, x(0)=2

myininaya (myininaya):

to use x'(0)=0 find x' \[x'=-c_1e^{-1t}-4c_2e^{-4t}\] \[0=-c_1-4c_2\] and to use x(0)=2 we have \[2=c_1+c_2+2\]

OpenStudy (anonymous):

could someone help expand on how to get the particular solution? you just set the left to be a constant A and solve x' and x''?

myininaya (myininaya):

i think you have to sort of guess what the form of particular solution will be like we could try x=A+Bt+Ct^2 and see what happens x'=0+B+2Ct x''=2C so plugin' these in we get 2C+5(B+2Ct)+4(A+Bt+Ct^2)=8 we need to get our t^2 together our t's together and our constants 4t^2*C+t*(10C+4B)+2C+5B+4A=8 but C=0 so B=0 then 4A=8=> A=2

OpenStudy (anonymous):

Since it is equal to a constant you can assume the particular part is a constant (or at least contains one as myininaya showed) As long as you don't guess sin(x) or cos(x) or anything. The method is called undetermined coefficients. This is a great website: http://tutorial.math.lamar.edu/Classes/DE/UndeterminedCoefficients.aspx

OpenStudy (anonymous):

how do we know C =0?

myininaya (myininaya):

there are no t^2 on the other side of the equation

OpenStudy (anonymous):

and B=0 because there are no t's either?

myininaya (myininaya):

right we have (10C+4B)t=0t 10C+4B=0 but we said first C=0 so we have 4B=0 so B=0

OpenStudy (anonymous):

and is this A =2 our t_p?

OpenStudy (anonymous):

There is a theorem in linear algebra that says: \[at^n+bt^{n-1}...=\alpha t^n+\beta t^{n-1}... \implies a=\alpha; b=\beta\] Though I gave a very "specific" by assuming it was a polynomial, it can be anything really.

myininaya (myininaya):

yes dof

OpenStudy (anonymous):

k, I think I got it; I appreciate all the time and help!

myininaya (myininaya):

np the only thing left is two solve your two equations above for c1 and c2

myininaya (myininaya):

but you can do that using algebra